The addition of a load of mass 2000kg to a car of mass 1200kg caused the spring of shock absorber to vibrate, if the spring force constant is 6.5 x 10⁴Nm
1
Expert's answer
2021-03-14T19:19:28-0400

We can find the period of vibration of the car as follows:


"T=2\pi\sqrt{\dfrac{m}{k}},""T=2\pi\sqrt{\dfrac{3200\ kg}{6.5\cdot10^4\ \dfrac{N}{m}}}=1.40\ s."